• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name boeglin

    Log entry time 16:02:10 on March15,2005

    Entry number 141590

    keyword=Shift Summary

    8:00 currently running 2676

    8:15 start run 2677
    8:40 beam lost due to RF problem. Move Hall to controlled access for work.

    9:30 E. Folts in Hall for work
    S. Spiegel in Hall for Work
    11:00 S. Spegel back out of the Hall
    11:45 E. Folts out of the Hall, Visitors back from Hall
    12:15 R. Feuerbach in Hall working on HRSL
    13:00 R. Feurbach back out of the Hall

    13:55 MCC asks to bring the Hall to Beam Permit
    14:35 CW beam in Hall

    runs 2678 - 2692 cosmic and test runs during beam studies.

    14:38 start run 2692 ->coda problem, restart. center the beam
    14:43 start run 2693 S2m TDC's are empty !!! -> investigate
    15:35 ask for a controlled access to fix TDC problem